# Information-theoretic description of a feedback-control Kuramoto model

### v1

@@ III.3 Viability Measure $\mathcal{V}$
Integrating over all coupling strengths, the viability, is shown in the bottom panels of Fig. 4. In both cases, we see a linear decrease in $\mathcal{V}$ with increasing $\eta$. The best linear fits to the curves (plotted as pink dashed lines), imply a constant viability per bit (VpB) in the agent/environment communication channel. For Poisson graphs VpB is found to be $0.171\pm 0.002$, while for scale-free graphs $0.209\pm 0.003$. Note that in the absence of noise, the presence of an agent that can accurately estimate the local mean field, has a net effect of lowering the global synchronization $\approx 20\%$ across the given range of coupling parameters. This appears to be independent of the topology, although the decrease in viability with noise is more pronounced in the scale-free case. This is likely due to the fact that the agent spends most of its time located in the hubs (given that the probability to visit a node $i$ in a random walk, is $p_{i}\propto k_{i}$), and according to Eq. (III.2) the computational timescale for hubs is much longer than lower-degree nodes. Furthermore, with increasing noise, the error in the estimation of the mean-field is enhanced due to the greater number of neighbors of a hub as compared to nodes in the Poisson graph.
@@
The linear dependence of the viability on the noise, the absence of any plateau, and a constant VpB is in stark contrast with the results seen in the forager model Sowinski et al. [2023]. There, the viability plateau—marked by a peak in the VpB—differentiated the region of semantic and syntactic information. Correlations above the threshold were irrelevant to the forager’s viability. In this setting, however, every bit of syntactic information is semantic. All correlations the agent has with the environment is essential to the viability of the combined agent-environment system.
@@ IV Discussion
The goals of this paper have been twofold. First we propose a general roadmap towards applying semantic information theory (SIT) theory to dynamical systems. Building on Kolchinsky and Wolpert [2018]; Sowinski et al. [2023] we articulate the steps required to cast a system into a form in which one can go beyond syntactic measures of information and consider instead how much of that information contributes to a systems self-maintenance (i.e. viability). Broadly, the procedure involves separating the system into agent-environment sub-systems; defining system dynamics; defining a measure of viability; identifying the state space of correlations between agent and environment such that information theoretic measures may be applied; defining a procedure to intervene in the dynamics; computing the semantic content and viability per bit from the ensemble of intervened trajectories. In this way a system and its dynamics can be analyzed as agent-environment correlations directly responsible for an agent’s effectiveness at maintaining itself, the environment, or both, far-from-equilibrium. As we have noted it is the the equilibrium state in question that provides the basis for a semantic interpretation of this information. Thus determination of such equilibria becomes an essential step in the applying SIT to a system. In some cases such equilibria may be obvious, such as the death of an agent Sowinski et al. [2023], however, as we have shown in this paper it is possible that the equilibria of interest, and hence proper viability functions, may require more effort to determine.
@@
Our second goal was to operationalize the road map by exploring a specific model to demonstrate how SIT can bring new insights to previously studied problems. Loosely motivated by biological models such as the neurodynamics of epilepsy and other biologically pathological states, we introduced an agent-based variant of the well-known Kuramoto model. Treating the original model on an arbitrary network as an environment, with an exogenously dependent coupling, an agent is introduced tasked with preventing/delaying the onset of synchronization when the environmental coupling exceeds the critical value. The modified system is a stochastic version of the original model where the effect of the agent is to function as a heat bath that deforms the distribution of frequencies for the oscillators, thus delaying the onset of synchronization. Articulating the resulting correlational structure for agent and environment (the oscillators), allows us to cast the model in in a novel computational (information theoretic) perspective, where the agent-environment dynamics can be thought of as analyzing a communication channel. Adding noise to this channel enables the calculation of the system’s viability as a function of the level of scrambling of the channel and the viability per bit. Unlike previous explorations of semantic information Sowinski et al. [2023], our results showed no semantic threshold. Instead we find a smooth linear decrease in viability with increased scrambling of the communication channel. The results indicate that in this case, syntactic and semantic information are indistinguishable; every bit of information in the communication channel between the agent and environment is essential for the system’s viability.
@@ Appendix A Simulation methodology
To acquire the data discussed in this paper we performed simulations written in MATLAB and run on the Bluehive cluster at the University of Rochester. For the environment we constructed $1024$ distinct graphs for each of the $256$ values of the coupling, $\sigma\in[0,1]$. The environments were initialized with random phases drawn uniformly, and random natural frequencies drawn from a Gaussian distribution with unit standard deviation. Using $\Delta t=1/16$, each environment was run for $1500$ timesteps to ensure the decay of any transients; data was then collected for the following $512$ timesteps.
With the introduction of the agent, $\alpha=0.5$ $\beta=0.01$, the number of distinct graphs was reduced to $32$, leaving the other numbers mentioned above the same. Furthermore, the experiments were now done for an additional $128$ values of the scrambling parameter, $\eta$. For simulations with and without agents, the underlying graphs were either Poisson or scale-free. The graphs were constructed via the configuration model Newman [2018]. The resulting graphs had $N=1000$ nodes and average degrees of $\langle k\rangle\approx 4$.
@@ Appendix B Derivation of Continuum Dynamics
The agent moves between nodes on a timescale $\tau$. During a time interval $\Delta t$, the agent makes $J=\Delta t/\tau=\beta N\Delta t$ jumps between nodes. Due to the topology of the environment the agent visits nodes with higher degree more often than those with lesser degree. The probability of visting a node $i$ is proportional to its degree, $p_{i}\propto k_{i}$ (with normalization $N\langle k\rangle$). Each visit results in a perturbation of strength $\alpha$, so the expected perturbation to that node is $\Delta\theta_{i}=\alpha p_{i}J=\alpha\kappa_{i}\beta k_{i}\Delta t/\langle k\rangle$, Since visits are rare, the perturbations are a Poisson arrival process. On timescales much larger than a single jump, the expected number of visits is high enough to model the Poisson process as a fixed rate with Gaussian fluctuations. Taken together, this means that we can think of the agent’s actions on each oscillator as approximated by a continuous perturbation of strength
@@
Note the presence of the $\sqrt{\Delta t}$ above, which motivates our quantification of this heuristic by the replacement of this term with a Wiener process.
@@ Appendix C Review of Fully Connected Model
In this appendix we review the calculation of the critical coupling in the original fully connected model. There we have that $\sigma=K/N$; the coupling needs to scale inversely with the number of nodes because in a fully-connected graph the number of connections, by definitions, scales with $N$. Our starting point is therefore Eq.  (4), the mean field equation of motion. We work under the simplifying assumption of the long-time, thermodynamic limit, $t\rightarrow\infty,N\rightarrow\infty$, so that fluctuations in the synchronization order parameter, $r$, vanish, allowing us to treat it as constant. In what follows we will derive a self-consistency constraint on the synchronization, from which everything else will follow.
Consider an oscillator with natural frequency $\omega$. If $|\omega|>Kr$, then $\dot{\theta}$ never vanishes, and the oscillator is unsynchronized. On the other hand if $|\omega|\leq Kr$, the oscillator synchronizes, approaching a phase of $\theta\rightarrow\phi+\sin^{-1}\omega/Kr$. We use these facts to break apart the sum in the definition of the synchronization into two sums, one over the synchronized, and the other over the unsynchronized, oscillators
$re^{i\phi}$ $=\frac{1}{N}\sum_{i}e^{i\theta_{i}}$
$=\frac{1}{N}\sum_{i|\omega_{i}>Kr}e^{i\theta_{i}}+\frac{1}{N}\sum_{i|\omega_{i}\leq Kr}e^{i\theta_{i}}$
The thermodynamic limit assumption allows us to disregard the first term - the unsynchronized oscillators will be distributed uniformly across phases, cancelling. For the second term, we use the long-time assumption to set the phases to their synchronized values noted above, resulting in
$r=\frac{1}{N}\sum_{i|\omega_{i}\leq Kr}e^{i\sin^{-1}\frac{\omega_{i}}{Kr}}$
The thermodynamic limit assumption here allows us to replace the sum with an appropriate integral over the frequency distribution, which when combined with some elementary trigonometric identities, results in
@@
Since we work with symmetric frequency distributions, the imaginary piece of the above vanishes. In what remains, we perform the substitution $\omega=Krx$, resulting in the self-consistency integral constraint
@@
Figure 5: Synchronization in the fully connected Kuramoto model with a Gaussian distribution of natural frequencies. The vertical line is at the critical coupling, $K_{c}=2\sqrt{2/\pi}$. The dashed curve is the second order approximation.
@@
where we have used the symmetry of $g$ as an even function. Plugging into the self-consistency constraint results in integrals easily recognizable as Euler beta functions, which massaged with several Euler gamma function identities result in
@@
The next order of the calculation,
$\frac{1}{K}$ $=\frac{1}{K_{c}}+\frac{\pi}{16}g^{\prime\prime}(0)K^{2}r^{2}$
@@
$r^{2}$ $=8\frac{g(0)}{g^{\prime\prime}(0)}\frac{K-K_{c}}{K^{3}}$ (32)

### v3

@@ III.3 Viability Measure $\mathcal{V}$
In both Poisson and scale-free graphs, we no longer see the emergence of a semantic threshold, with a smooth linear dependence of the viability curves with sensor fidelity. The best linear fits to the curves (plotted as pink dashed lines), imply a constant viability per bit (VpB) in the agent/environment communication channel. For Poisson graphs VpB is found to be $0.886\pm 0.01$, while for scale-free graphs $0.841\pm 0.01$. Note that in the absence of noise, the presence of an agent that can accurately estimate the local mean field, has a net effect of lowering the global synchronization $\approx 20\%$ across the given range of coupling parameters. This appears to be independent of the topology, although the decrease in viability with noise is more pronounced in the scale-free case. This is likely due to the fact that the agent spends most of its time located in the hubs (given that the probability to visit a node $i$ in a random walk, is $p_{i}\propto k_{i}$), and according to Eq. (III.2) the computational timescale for hubs is much longer than lower-degree nodes.
@@
The absence of any plateau, and a constant VpB in the latter two graphs is in stark contrast with the results seen in the fully-connected one and in prior models Sowinski et al. [2023]. There, the viability plateau—marked by a peak in the VpB—differentiated the region of semantic and syntactic information. Correlations above the threshold were irrelevant to the forager’s viability. In these cases, however, every bit of syntactic information is semantic. All correlations the agent has with the environment are essential to the viability of the combined agent-environment system. The fact that the threshold appears in the fully connected case, whereas the qualitative behavior of the viability is indistinguishable in the two sparser graphs, suggests that the key parameter in determining the presence/absence of the threshold is not $\langle k^{2}\rangle$, but the density of the graph $\langle k\rangle/N$.
@@ IV Discussion
The goals of this paper have been twofold. First we propose a general road map towards applying semantic information theory (SIT) theory to dynamical systems. Building on Kolchinsky and Wolpert [2018]; Sowinski et al. [2023] we articulate the steps required to cast a system into a form in which one can go beyond syntactic measures of information and consider instead how much of that information contributes to a systems self-maintenance (i.e. viability). Broadly, the procedure involves separating the system into agent-environment sub-systems; coarse-graining the DoF and identifying which correlations to follow; identifying an intrinsic viability measure; intervening with a scrambling protocol on the correlations, and measuring the change in viability. In this way a system and its dynamics can be analyzed as agent-environment correlations directly responsible for an agent’s effectiveness at maintaining itself, the environment, or both, far-from-equilibrium. As we have noted it is the equilibrium state in question that provides the basis for a semantic interpretation of this information. Thus determination of such equilibria becomes an essential step in the applying SIT to a system. In some cases such equilibria may be obvious, such as the death of an agent Sowinski et al. [2023], however, as we have shown in this paper it is possible that the equilibria of interest, and hence proper viability functions, may require more effort to determine.
@@
Our second goal was to operationalize the road map by exploring a specific model to demonstrate how SIT can bring new insights to previously studied problems. Loosely motivated by biological models such as the neurodynamics of epilepsy and other biologically pathological states, we introduced an agent-based variant of the well-known Kuramoto model. Treating the original model on an arbitrary network as an environment, with an exogenously dependent coupling, an agent is introduced tasked with preventing/delaying the onset of synchronization when the environmental coupling exceeds the critical value. The modified system is a stochastic version of the original model where the effect of the agent is to function as a heat bath that deforms the distribution of frequencies for the oscillators, thus delaying the onset of synchronization. Articulating the resulting correlational structure for agent and environment (the oscillators), allows us to cast the model from a computational (information theoretic) perspective, where the agent-environment dynamics can be analyzed as a communication channel. Adding noise to this channel enables the computation of the system’s viability as a function of the level of scrambling of the channel and the viability per bit. As in previous explorations of semantic information Sowinski et al. [2023], our results showed a semantic threshold in fully-connected graphs. Both Poisson and scale-free graphs, however, lacked this feature. The results indicate that semantic information is affected by network topology, with the relevant parameter being the density of the graph as opposed to the degree distribution. For sparser graphs, every bit of information in the communication channel between the agent and environment is essential for the system’s viability, whereas dense graphs display a semantic threshold.
@@
V Acknowledgements
The authors thank Artemy Kolchinsky for his comments on the manuscript, and the Center for Integrated Research Computing (CIRC) at the University of Rochester for providing computational resources and technical support. A heartfelt thanks to the thorough readings and comments made by the anonymous reviewers for which we credit our discovery of the semantic threshold in the fully-connected model. This project was partly made possible through the support of Grant 62417 from the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.
@@ Appendix A Simulation methodology
To acquire the data discussed in this paper we performed simulations written in MATLAB and run on the Bluehive cluster at the University of Rochester. For simulations with and without agents, the underlying graphs were either fully-connected, Poisson, or scale-free. The latter were constructed via the configuration model Newman [2018], resulting in graphs with $N\approx 1000$ nodes and average degree of $\langle k\rangle\approx 4$.
For the Poisson and scale-free environments we constructed $1024$ distinct graphs for each of the $256$ values of the coupling, $\sigma\in[0,1]$. The same was done for the full-connected environments, except that the coupling was in the range $K\in[0,4]$. The environments were initialized with random phases drawn uniformly, and random natural frequencies drawn from a Gaussian distribution with unit standard deviation. Using $\Delta t=1/16$, each environment was run for $1500$ timesteps to ensure the decay of any transients; data was then collected for the following $512$ timesteps and averaged to give the synchronization of each environment.
With the introduction of the agent, $\alpha=0.5$ $\beta=0.16$, the number of distinct Poisson and scale-free graphs was reduced to $32$, leaving the other numbers mentioned above the same. Furthermore, the experiments were now done for an additional $128$ values of the scrambling parameter, $\eta$, equally spaced in the range $[0,1]$.
@@ Appendix B Derivation of Continuum Dynamics
The agent moves between nodes on a timescale $\tau$. During a time interval $\Delta t$, the agent makes $J=\Delta t/\tau=\beta N\Delta t$ jumps between nodes. Due to the topology of the environment the agent visits nodes with higher degree more often than those with lesser degree. The probability of visiting a node $i$ is proportional to its degree, $p_{i}\propto k_{i}$ (with normalization $N\langle k\rangle$). Each visit results in a perturbation of strength $\alpha$, so the expected perturbation to that node is $\Delta\theta_{i}=\alpha p_{i}J=\alpha\kappa_{i}\beta k_{i}\Delta t/\langle k\rangle$, Since visits are rare, the perturbations are a Poisson arrival process. On timescales much larger than a single jump, the expected number of visits is high enough to model the Poisson process as a fixed rate with Gaussian fluctuations. Taken together, this means that we can think of the agent’s actions on each oscillator as approximated by a continuous perturbation of strength
@@
Note the presence of the $\sqrt{\Delta t}$ above, which motivates our quantification of this heuristic by the replacement of this term with a Wiener process. To see how well the continuum model works we plot it against the agent based model in figure 5.
Figure 5: In the top panel we have a comparison of the order parameter in the the agent-based model (solid curve) and the stochastic differential equation eq.III.1.4 (dashed curve). The bottom panel shows the absolute error between the two. Note that the y-axis is logarithmic in order to exaggerate the underestimation of the SDE.
@@ Appendix C Review of Fully Connected Model
In this appendix we review the calculation of the critical coupling in the original fully connected model. There we have that $\sigma=K/N$; the coupling needs to scale inversely with the number of nodes because in a fully-connected graph the number of connections, by definition, scales with $N$. Our starting point is therefore Eq.  (4), the mean field equation of motion. We work under the simplifying assumption of the long-time, thermodynamic limit, $t\rightarrow\infty,N\rightarrow\infty$, so that fluctuations in the synchronization order parameter, $r$, vanish, allowing us to treat it as constant. In what follows we will derive a self-consistency constraint on the synchronization, from which everything else will follow.
Consider an oscillator with natural frequency $\omega$. If $|\omega|>Kr$, then $\dot{\theta}$ never vanishes, and the oscillator is unsynchronized. On the other hand if $|\omega|\leq Kr$, the oscillator synchronizes, approaching a phase of $\theta\rightarrow\phi+\sin^{-1}\omega/Kr$. We use these facts to divide the set of oscillator indices, $\mathcal{I}$, into synchronized, $\mathcal{I}_{s}=\{i:|\omega_{i}|<Kr\}$, and unsynchronized, $\mathcal{I}_{u}=\mathcal{I}\backslash\mathcal{I}_{s}$, subsets. The sum in the definition of the order parameter becomes
$re^{i\phi}$ $=\frac{1}{N}\sum_{i\in\mathcal{I}}e^{i\theta_{i}}$
$=\frac{1}{N}\sum_{i\in\mathcal{I}_{u}}e^{i\theta_{i}}+\frac{1}{N}\sum_{i\in\mathcal{I}_{s}}e^{i\theta_{i}}$
The thermodynamic limit assumption allows us to disregard the first term - the unsynchronized oscillators will be distributed uniformly across phases, cancelling. For the second term, we use the long-time assumption to set the phases to their synchronized values noted above, getting
$r=\frac{1}{N}\sum_{i\in\mathcal{I}_{s}}e^{i\sin^{-1}\frac{\omega_{i}}{Kr}}$
Then we insert an integral over frequencies containing the delta function $\delta(\omega-\omega_{i})$. The sum gives a density of states, which in this case is the frequency distribution and a Jacobian factor in going from oscillator index to oscillator frequency, resulting in
@@
Since we work with symmetric frequency distributions, the imaginary piece of the above vanishes. In what remains, we perform the substitution $\omega=Krx$, to get the self-consistency constraint
@@
Figure 6: Synchronization in the fully connected Kuramoto model with a Gaussian distribution of natural frequencies. The vertical dashed-dot line is at the critical coupling, $K_{c}=2\sqrt{2/\pi}$. The dashed horizontal line is the Rayleigh distribution mean. The light grey region is data from our simulations, with the mean a solid curve going through it. The black curve is the analytical solution eq. 25. The dashed blue curve is the second order approximation to that solution.
@@
where the even symmetry of $g$ has been used. Plugging into the self-consistency constraint results in integrals easily recognizable as Euler beta functions. Rewriting this in terms of Euler gamma function identities results in
@@
If the distribution is monotonically decreasing for $\omega>0$, and does not have vanishing curvature at $\omega=0$, then we can go to second order to extract the behavior near the critical coupling. Analysis of
$\frac{1}{K}$ $=\frac{1}{K_{c}}-\frac{\pi}{16}|g^{\prime\prime}(0)|K^{2}r^{2}$
@@
$r^{2}$ $=8\frac{g(0)}{|g^{\prime\prime}(0)|}\frac{K-K_{c}}{K^{3}}$ (29)
@@
Appendix D Critical Coupling with an Agent
Here we present several methods for estimating the critical coupling in the presence of an agent. The first of utilizes a Gaussian approximation for the shape of the effective distribution, whereas the next two use a parameterized ansatz for the effective distribution.
D.1 Gaussian Approximation
Here we approximate the effective distribution as a Gaussian, showing that the increase in variance when $g(0)$ is replaced with $g^{A}(0)$ in equation 28 results in a $K_{c}^{A}>K_{c}$. As described at the end of section III.1.4, the frequency distribution is first cut at $\omega=0$, and the positive (negative) half shifted by positive (negative)$\alpha\beta$. When this is done for a standardized Gaussian, a quick calculation finds the mean unchanged, and the variance now $1+2\sqrt{2/\pi}\alpha\beta+\alpha^{2}\beta^{2}$. Next, the convolution with a heat kernel of variance $\alpha\beta$, results in the variance of the effective distribution,
$\langle\omega^{2}\rangle_{A}$ $=1+(1+2\sqrt{\frac{2}{\pi}})\alpha\beta+\alpha^{2}\beta^{2}$ (30)
$\geq 1.$
We next assume the effective distribution is also a Gaussian, with the above variance. The peak of the effective distribution related to the original is
$g^{A}(0)=\frac{g(0)}{\langle\omega^{2}\rangle_{A}^{1/2}}$ (31)
The critical coupling is consequently shifted according to,
$K_{c}\rightarrow K^{A}_{c}$ $=\langle\omega^{2}\rangle_{A}^{1/2}K_{c}.$ (32)
In our simulations, $\alpha\beta=0.08$, this corresponds to $K^{A}_{c}\approx 1.75829$. This however is an underestimate of the true critical coupling.
D.2 Parameterized Effective Distribution
In order for an improved estimate, we next calculate the critical coupling from the parameterized effective distribution rather than approximating it as a Gaussian. The initial cut in the distribution and shift of frequencies results in
$g(\omega)\mapsto\tilde{g}(\omega)$ $=\Theta(-\alpha\beta-\omega)g(\omega+\alpha\beta)$
$\hskip 10.0pt+\Theta(\omega-\alpha\beta)g(\omega-\alpha\beta),$ (33)
where $\Theta$ is the Heaviside step function. It is easily seen that the first term is the negative part of the distribution shifted to the left, and the second is the positive part shifted to the right. Convolving with the heat kernel, $h_{\alpha\beta}$, results in
[…]
$=\frac{1}{2\sqrt{2\pi}\sqrt{1+\alpha\beta}}\left(e^{-\frac{1}{2}\frac{(\omega+\alpha\beta)^{2}}{1+\alpha\beta}}\text{erfc}\left[\frac{\omega+\alpha\beta}{\sqrt{2\alpha\beta(1+\alpha\beta)}}\right]+e^{-\frac{1}{2}\frac{(\omega-\alpha\beta)^{2}}{1+\alpha\beta}}\text{erfc}\left[\frac{-\omega+\alpha\beta}{\sqrt{2\alpha\beta(1+\alpha\beta)}}\right]\right),$ (36)
where erfc$(x)$ is the complementary error function. This suggests a bimodal form for $g^{A}(\omega)$ constructed of two overlapping Gaussian distributions; the smaller $\alpha\beta$ is, the greater the overlap, giving the appearance of a Gaussian with a dimple at its center. Several of these effective distributions are plotted in the upper left panel of figure 7. The solid black curve represents simulation data ($\alpha\beta=0.08$), and the thin blue curves are shown for a range $\alpha\beta$ above and below 0.08. For comparison the original Gaussian distribution, $g(\omega)$, is plotted as a dashed line.
Figure 7: In the top left panel we have the original Gaussian frequency distribution, $g(\omega)$, plotted with a dashed line. The thick black curve is the effective distribution, $g^{A}(\omega)$, while the thin blue lines are representative of smaller and higher $\alpha\beta$. The top right panel is the behavior of the critical coupling as a function of $\alpha\beta$; the full non-linear form shown in black, and the linearized form in pink. The dashed line is the corresponding estimate for $K_{c}^{A}$. The bottom panel shows the full analysis of the consistency constraint. Details are in the text.
Plugging in $\omega=0$ into the effective distribution, and then inserting the result into equation 28, shows how $\alpha\beta$ parameterizes the change in the critical coupling
$K_{c}^{A}$ $=\sqrt{1+\alpha\beta}\frac{e^{\frac{1}{2}\frac{(\alpha\beta)^{2}}{1+\alpha\beta}}}{\text{erfc}\left[\sqrt{\frac{\alpha\beta}{2(1+\alpha\beta)}}\right]}K_{c}$ (37)
$\approx(1+\sqrt{\frac{2}{\pi}\alpha\beta}+\frac{\pi+4}{2\pi}\alpha\beta+\cdots)K_{c}.$ (38)
In the second line we have expanded for $\alpha\beta\ll 1$; we see that the critical coupling is a Taylor series in $\sqrt{\alpha\beta}$. Furthermore, computing the second derivative of the effective potential gives a critical exponent of $1/2$ for the behavior of the new critical coupling; in the Gaussian approximation this critical exponent is $1$. This behavior of $K_{c}^{A}/K_{c}$ is plotted in the upper right panel of figure 7. The solid black is the full non-linear behavior, while the straight pink line is the linearized behavior (quadratic in $\sqrt{\alpha\beta}$). The dashed line is at $\alpha\beta=0.08$ giving us a second estimate for the critical coupling, $K_{c}^{A}\approx 2.11751$.
This estimate turns out to be a bit high. Getting an even better estimate of the critical point is difficult analytically, as the consistency constraint, to the best of our knowledge, cannot be integrated by hand. Numerical integration, on the other hand, can be quickly implemented to find $r=r(K)$, which is plotted in the bottom panel of figure 7 with the solid black curve. The order parameter is multi-valued in the range $K\in[1.9887,2.118]$, where the right boundary coincides with the previous estimate. Plotting this alongside the simulation data (the grey region) shows us the estimate is high and that the left end of the range is a better estimate of the critical coupling. Both of these are plotted as the vertical dashed lines in the panel. The dashed pink curve is the expected value of the synchronization, which matches very well with numerical simulations. In the thermodynamic limit the synchronization curve has a discontinuity at $K_{c}^{A}\approx 1.9887$, and drops to 00 at smaller values of the coupling.
The plot of the consistency constraint of the effective distribution in figure 7 matches the major features of the data. This gives us reasonable confidence that the above analysis using the method of an effective distribution is a valid way of describing the agent-environment system. That said, the match is not perfect, as one can see from the slight underestimation of the data in the figure. Since the ensemble size is large, $N=1024$, we do not think that this is due to not being at the thermodynamic limit, allowing us to conjecture that there is yet some refinement to this method which we leave for future work.